# Modeling Binary Lenses and Sources with the BAGLE Python Package

### v1

@@ 4.3 Full Keplerian Solutions
4.4 Lensing a binary source
@@
Each source’s intrinsic flux is magnified by its specific amplification factors. The total amplification for each source is
@@
We can define a total amplification for the system using
@@
$f_{obs}=(A_{p}F_{p}+A_{s}F_{s})\left(1+\frac{1-b_{sff}}{b_{sff}}\right).$ (17)
where $b_{sff}$ is the ratio of the source flux to the total flux of the source, neighbors, and the lens.
The lensed astrometry, or the image centroid, is then simply a flux-weighted combination of the lensed astrometry from the two sources.
5 Point Sources and Binary Lens (PSBL)
In this section, we begin by discussing the static lens approximation in §5.1, followed by the linear and accelerated orbital approximations in §5.2. In §5.3, the full Keplerian solutions are presented. Lastly, we discuss the microlensing equations for binary lenses in §5.4.
5.1 Static Approximation
@@
For binary lenses, $t_{0,\sun}=t_{geom,0,\sun}$ in static, linear and accelerated orbital approximations. $t_{geom,0,\sun}$ represents the time of closest approach between the source and the geometric midpoint of the binary lenses. In the static approximation only, the angular separation between the primary and secondary lens is fixed at all times; both lenses move with the same proper motion vector.
5.2 Linear and Accelerated Orbital Approximations
Figure 5: Source and lens trajectories for linear (upper panel) and accelerated (lower panel) approximations of orbital motion of binary lenses. The solid black line is the primary lens, the solid purple line is the secondary lens, and the solid yellow lines are the image positions. The green line is the flux-weighted average of the lensed source position, as observed on the sky. Note that $mag_{S}$ = 16. In both panels, the source is stationary; the primary lens has a proper motion of $\boldsymbol{\mu}_{\boldsymbol{L},\sun}$ = [-3.76 mas yr-1, -3.76 mas yr-1]; the secondary lens has a proper motion of $\boldsymbol{\mu}_{\boldsymbol{L_{s}},\sun}$ = [-2.76 mas yr-1, -2.76 mas yr-1]. For our model with acceleration, we provide the following input for $\mathbb{a}_{\boldsymbol{L\mathrm{rel}},\sun}$ = [1 mas yr-1, -1 mas yr-1]. In both cases, the Einstein time is $t_{E,\sun}$=412 days and $u_{0,\sun}$ = 0.5. Note that this scenario does not involve a caustic crossing and thus produces only 3 images.
The linear orbital approximations for binary lenses follow the same logic as binary sources in §4.2. Therefore, after accounting for parallactic motion (given by $\pi_{L}\boldsymbol{P}(t,\alpha,\delta)$ where $\pi_{L}$ is the maximum parallax amplitude of the lens, $1/d_{L}$), the positions of the primary and secondary lenses are given by:
@@ 5.2 Linear and Accelerated Orbital Approximations
5.3 Full Keplerian Solutions
Figure 6: Source and lens trajectories for a simulated microlensing event with a binary lens. The solid black line is the primary lens, the solid purple line is the secondary lens, the orange point is the source position, and the solid yellow lines are the image positions. The green line is the flux-weighted average of the lensed source position, as observed on the sky. Note that $mag_{S}$ = 16. We use the following orbital parameters: $\omega_{pri}=30\degree$, $\Omega_{sec}=10\degree$, $\textit{i}=90\degree$, $e=0.6$, $\textit{P}=1054.41$ days, and an angular separation of 55 mas between the two lenses at $t_{com,0,\sun}$. We present our simulation over $t_{E}=412.02$ days. The lenses have a mass of $m_{L,p}=$ 10 $M_{\odot}$ and $m_{L,s}=$ 5 $M_{\odot}$. The source is stationary.
All binary-lens models with Keplerian orbital motion (either circular or elliptical) input the following new Keplerian elements: $\omega_{pri}$, $\Omega_{sec}$, i, e, $t_{p}$, and $a$, where $a$ is the magnitude of the semi-major axis in angular units (mas) at $t_{p}$. We can convert $a$ to the semi-major axis in units of AU ($a_{AU}$) by using $a$ and the distance to the lens ($d_{L}$) in units of pc, i.e., $a_{AU}=a\times d_{L}$. In these models, $\boldsymbol{\mu}_{\boldsymbol{L},\sun}$ is treated as the proper motion of the binary lens system’s center of mass instead of the geometric midpoint of the binary lens. As such, we change the notation from $\boldsymbol{\mu}_{\boldsymbol{L},\sun}$ to $\boldsymbol{\mu}_{\boldsymbol{L,com},\sun}$ throughout this section. Furthermore, we amend the definition of $t_{0,\sun}$ for such binary-lens models, i.e, $t_{0,\sun}=t_{com,0,\sun}$, where $t_{com,0,\sun}$ is the time of closest approach between the lens center of mass and the source. The initial position of the lens system’s center of mass at $t_{com,0,\sun}$ is input as $\boldsymbol{X}_{\boldsymbol{L_{com},0},\sun}$.
@@ 5.3 Full Keplerian Solutions
Using all eight Keplerian elements, the motion of the primary and secondary lenses around their center ($\boldsymbol{X(t)}_{\boldsymbol{p},\boldsymbol{com},\sun}$ for primary and $\boldsymbol{X(t)}_{\boldsymbol{p},\boldsymbol{com},\sun}$ for secondary) of mass at rest can be calculated. Once, the proper motion of the center of mass and the parallactic motion are taken into account, the primary and secondary lens trajectories are given by:
@@
We can simulate binary lens astrometry trajectories with Keplerian solutions implemented in BAGLE, presented in Figure 6. The code necessary to simulate the astrometric trajectories is displayed below. We use the PSBL parameterization PSBL_PhotAstrom_noPar_EllOrbs_Param1, while noting that BAGLE has multiple alternative parameterizations.
@@
3#Create a Keplerian orbital model with photometric and astrometric parameters
4
5psblorbits = model.PSBL_PhotAstrom_noPar_EllOrbs_Param1(
@@
7 beta, muL_E, muL_N, omega, big_omega, i, e, tp, a, muS_E, muS_N, dL, dS,
8 b_sff, mag_src,dmag_Lp_Ls,
9 raL=None, decL=None, root_tol=1e-8
@@
11# Get resolved astrometry for lenses
12 lens1, lens2 = psbl.get_resolved_lens_astrometry(t)
13
14# Get unlensed source trajectory
15source_unlensed = psbl.get_astrometry_unlensed(t)
@@
17# Get resolved lensed images
18images_resolved = psbl.get_resolved_astrometry(t, image_arr=img, amp_arr=amp)
5.4 Binary Lens Equation
@@
The complex conjugate of Equation  27 is:
@@
Differentiating Eqns. 27 and 32 gives
@@
We can then find the amplification of the source by the lens as:
@@
There are places in which $|J|\rightarrow 0$. This corresponds to infinite amplification (if the source were a point). The curves in the lens plane where this is true are “critical curves” and the corresponding curves in the source plane are called “caustics.” In the maps of magnification below (see §11), we show the source plane in which we can see caustics. We also show the caustics in Figures 13 and 14. When the source passes over a caustic, this is known as a “caustic crossing” and changes the number of lensed images of that source from 3 (when the source is outside of the caustic) to 5 (when it is inside the caustic).
@@
In order to avoid working with an equation with a mix of complex numbers and their conjugates, it is standard to plug Eqn. 32 into Eqn. 27 and simplify (see H. J. Witt, 1990; H. J. Witt & S. Mao, 1995). This yields a fifth-order polynomial known as the “lens polynomial.” This typically yields the same solutions as the “lens equations.”
6 Binary Sources and Binary Lenses (BSBL)
Figure 8: Source and lens trajectories for a simulated binary source, binary lens microlensing event with involving circular orbits. The orange line is the primary source, the pink line is the secondary source, the black line is the primary lens, the purple line is the secondary lens, and the yellow lines are the image positions. The green line is the flux-weighted average of lensed source positions, as observed on the sky. Note that $mag_{S,pri}$ = 16 and $mag_{S,sec}$ = 17. The simulation was run with a $t_{E}=231.15$ days with the following orbital parameters for the lens: $\omega_{pri}=30\degree$, $\Omega_{sec}=10\degree$, $\textit{i}=90\degree$, $e=0.2$, $\textit{P}=2722.46$ days, and the following orbital parmaeters for the source $\omega_{pri}=30\degree$, $\Omega_{sec}=10\degree$, $\textit{i}=90\degree$, $e=0.4$, $\textit{P}=6000$ days. The lenses have a mass of $m_{L,p}=$ 10 $M_{\odot}$ and $m_{L,s}=$ 8 $M_{\odot}$.
BAGLE can also simulate events with both binary sources and binary lenses. The orbital and lensing equations necessary to simulate a binary source and binary lens event in BAGLE are individually handled. This means that the equations necessary to simulate binary sources are from §4, and the equations necessary to simulate binary lenses are from §5.
@@ 6 Binary Sources and Binary Lenses (BSBL)
We present our astrometric simulations in Figure 8. BSBL_PhotAstrom_noPar_EllOrbs_Param2 is the binary source and binary lens model from BAGLE that is used to simulate an orbital solution for such a microlensing event. The BSBL model assumes that both - the source and the lens - display orbital motion. The figure includes the observed (unresolved) source position on the sky, i.e., the flux-weighted average of lensed source positions.
7 Validation of Models
Figure 9: Comparison of a simulated PSBL event with parallax between VBMicrolensing, MulensModel and BAGLE. The microlensing parameters in the SSB lens-frame are: $t_{0,\sun}=60478.49$ MJD, $u_{0,\sun}=1$, $t_{E,\sun}=100.4$ days, $q=0.3$, $\pi_{E}=[0.3,0.2]$. The amplification (top) and residuals with respect to the BAGLE model (bottom) are shown over time.
Figure 10: Comparison of a simulated BSPL event with parallax between VBMicrolensing, MulensModel and BAGLE. The microlensing parameters in the SSB lens-frame are: $t_{prim,0,\sun}=60478.49$ MJD, $t_{sec,0,\sun}=60476.09$ MJD, $u_{p,0,\sun}=1$, $u_{s,0,\sun}=1.22$, $t_{E,\sun}=100.4$ days, $q_{flux,s}=0.16$, $\pi_{E}=[0.3,0.2]$. The amplification (top) and residuals with respect to the BAGLE model (bottom) are shown over time.
@@ 7 Validation of Models
In this section, we compare BAGLE with other contemporary microlensing models (VBMicrolensing and MulensModel) against simulated PSBL and BSPL events with parallax, but without orbital motion. A comparison with the inclusion of orbital motion and a comparison with BSBL is reserved for the future.
@@
We begin by creating a BAGLE model using the parameterization PSBL_Phot_Par_Param1. PSBL_Phot_Par_Param1 inputs quantities with reference to the geometric midpoint of the lens system, and calculates Earth’s position relative to the Solar System Barycenter over time. Event parameters in BAGLE are: $t_{0,\sun}=60478.49$ MJD, $u_{0,\sun}=1$, $t_{E,\sun}=100.4$ days, $q=0.3$, $\pi_{E}=[0.3,0.2]$. The event is located in the Galactic Bulge. The relative angle between the binary lens system and the $\mu_{\mathrm{rel},\sun}$ directional vector is $\phi=125\degree$. The two lenses have a separation of $\boldsymbol{s_{L}}(t_{0,\sun})=0.8$. In contrast to BAGLE, other packages prefer a geo-projected frame of reference (as described in J. Lu et al. (submitted)). The resulting geo-projected parameters are $t_{0,\earth_{r}}=60464.64$ MJD, $u_{0,\earth_{r}}=0.98$, $t_{E,\sun}=65.06$ days, $q=0.3$, $\pi_{E}=[-0.34,-0.13]$. After converting to the geo-projected frame of reference, an additional transformation must be applied to $\boldsymbol{u}_{\earth_{r}}$ and $t_{0,\earth_{r}}$ in order to switch from the geometric midpoint of the lens to the center of mass. The final values input to VBMicrolensing and MulensModel are $t_{0,\earth_{r}}=60460.41$ and $\boldsymbol{u}_{\earth_{r}}=0.78$. The comparison between VBMicrolensing, MulensModel and BAGLE for PSBL events is shown in Figure 9. pyLIMA was not compared because it uses VBMicrolensing as a backend. Differences of 10-2 are apparent between BAGLE and VBMicrolensing. VBMicrolensing assumes the source has a finite size and uses contour integration, whereas BAGLE uses the root solver. We believe the differences arise due to this. On the contrary, difference between MulensModel and BAGLE are incredibly small by several orders of magnitude (10-14) since both codes use the root solver by default.
@@
Next, a BSPL event pointing towards the Galactic Bulge was compared between VBMicrolensing, MulensModel and BAGLE. The parameterization BSPL_Phot_Par_Param1 was used to create a BAGLE model with the following microlensing parameters in the SSB reference frame: $t_{prim,0,\sun}=60478.49$ MJD, $t_{sec,0,\sun}=60476.09$ MJD, $u_{p,0,\sun}=1$, $u_{s,0,\sun}=1.25$, $t_{E,\sun}=100.4$ days, $q_{flux,s}=0.16$, $\pi_{E}=[0.3,0.2]$; where $u_{s,0,\sun}$ is the closest approach between the secondary source and lens, $t_{sec,0,\sun}$ is the time of this closest approach, and $q_{flux,s}$ is the ratio between the secondary source’s flux and the primary source’s flux. In the geo-projected frame, $t_{prim,0,\earth_{r}}=60464.64$ MJD, $t_{sec,0,\earth_{r}}=60461.01$ MJD, $u_{p,0,\earth_{r}}=0.98$, $u_{s,0,\earth_{r}}=1.22$, $t_{E,\earth_{r}}=65.06$ days, $\boldsymbol{\pi}_{E,\earth_{r}}=[-0.34,-0.13]$. The comparison between different BSPL models is presented in Figure 10. Differences of 10-4 are apparent between BAGLE and VBMicrolensing, likely due to differences in the parallax implementations. On BAGLE and MulensModel have a much smaller difference around 10-15. This level of difference is consistent with the point-source, point-lens comparison (with parallax) presented in J. Lu et al. (submitted).
@@
Next, we compare the computational run time of BAGLE  with VBMicrolensing and MulensModel. These runtimes are based on the time taken to model the events presented in Figures 9 and 10. Runtimes are calculated for models by generating mock data for an event with $t_{E}=100$ days sampling 2000 time steps over 5.5 years. The time tests are repeated 100 times and the mean and standard deviation are recorded for test. Tests were performed with BAGLE, VBMicrolensing v5.3.3, and MulensModel v3.3.1. A 2021 iMax equipped with a M1 processor using a python 3.11 environment and the ipython kernel was used to conduct these tests. A summary of model-generation runtimes is presented in Table 1 and 2 for PSBL and BSPL events, respectively.
@@
The pre-instantiated runtime is applicable to running many different sets of model parameters for an event at the same sky location (i.e. model fitting). On the other hand, the full runtime is applicable to large-scale simulations of multiple events located at different sky coordinates. For pre-instantiated runtimes, we find that all packages are quick for BSPL events. BAGLE is the quickest for BSPL model generations when calculating the full runtime. However, BAGLE is the slowest for PSBL model generation (pre-instantiated and full). This is due to BAGLE’s reliance on a polynomial root solver to calculate the images and amplification during a PSBL event. The polynomial root solver is used to find the roots of a fifth-order complex polynomial for each source position and epoch in BAGLE. The root solver enforces higher numerical precision and ensures more stable solutions, including near caustics. However, it is more computationally intensive than the contour-based or grid-based ray-shooting methods used by other codes. Future work involves implementing JAX to improve the efficiency of the root solver.

### v3

@@ IV.3 Full Keplerian Solutions
IV.4 Lensing a binary source
@@
Each source’s intrinsic flux is magnified by its specific magnification factors. The total magnification for each source is
@@
We can define the effective magnification for the system using
@@
$f_{obs}=(A_{p}F_{p}+A_{s}F_{s})+(F_{p}+F_{s})\left(\frac{1-b_{sff}}{b_{sff}}\right).$ (16)
where $b_{sff}$ is the ratio of the total source flux to the total flux of the source, neighbors, and the lens.
The image centroid is then simply a flux-weighted combination of the lensed astrometry from the two sources.
V Point Sources and Binary Lens (PSBL)
In this section, we begin by discussing the static lens approximation in §V.1, followed by the linear and accelerated orbital approximations in §V.2. In §V.3, the full Keplerian solutions are presented. Lastly, we discuss the binary lens equations in §V.4.
V.1 Static Approximation
@@
For binary lenses, $t_{0,\sun}=t_{geom,0,\sun}$ in static, linear and accelerated orbital approximations. $t_{geom,0,\sun}$ represents the time of closest approach between the source and the geometric midpoint of the binary lenses. In the static approximation only, the angular separation between the primary and secondary lens is fixed at all times.
V.2 Linear and Accelerated Orbital Approximations
Figure 5: Source and lens trajectories for linear (upper panel) and accelerated (lower panel) approximations of orbital motion of binary lenses. The solid black line is the primary lens, the solid purple line is the secondary lens, and the solid yellow lines are the image positions. The green line is the flux-weighted average of the lensed source position, as observed on the sky. Note that $mag_{S}$ = 16. In both panels, the source is stationary; the primary lens has a proper motion of $\boldsymbol{\mu}_{\boldsymbol{L},\sun}$ = [-3.76 mas yr-1, -3.76 mas yr-1]; the secondary lens has a proper motion of $\boldsymbol{\mu}_{\boldsymbol{L_{s}},\sun}$ = [-2.76 mas yr-1, -2.76 mas yr-1]. For our model with acceleration, we provide the following input for $\mathbb{a}_{\boldsymbol{L\mathrm{rel}},\sun}$ = [1 mas yr-1, -1 mas yr-1]. In both cases, the Einstein time is $t_{E,\sun}$=412 days, the Einstein radius is 66 mas
and $u_{0,\sun}$ = 0.5. Note that this scenario does not involve a caustic crossing and thus produces only 3 images.
The linear orbital approximations for binary lenses follow the same logic as binary sources in §IV.2. Therefore, after accounting for parallactic motion (given by $\pi_{L}\boldsymbol{P}(t,\alpha,\delta)$ where $\pi_{L}$ is the parallax amplitude of the lens, $1/d_{L}$), the positions of the primary and secondary lenses are given by:
@@ V.2 Linear and Accelerated Orbital Approximations
V.3 Full Keplerian Solutions
Figure 6: Source and lens trajectories for a simulated microlensing event with a binary lens. The solid black line is the primary lens, the solid purple line is the secondary lens, the orange point is the source position, and the solid yellow lines are the image positions. The green line is the flux-weighted average of the lensed source position, as observed on the sky. Note that $mag_{S}$ = 16. The Einstein radius is $4.5$ mas
. We use the following orbital parameters: $\omega_{pri}=30\degree$, $\Omega_{sec}=10\degree$, $\textit{i}=90\degree$, $e=0.6$, $\textit{P}=1054.41$ days, and an angular separation of $5$ mas between the two lenses at $t_{com,0,\sun}$. We present our simulation over $t_{E}=412.02$ days. The lenses have a mass of $m_{L,p}=$ 10 $M_{\odot}$ and $m_{L,s}=$ 5 $M_{\odot}$. The source is stationary.
All binary-lens models with Keplerian orbital motion (either circular or elliptical) input the following new Keplerian elements: $\omega_{pri}$ (set to 0 for circular orbits), $\Omega_{sec}$, i, e, $t_{p}$, and $a$, where we define $a=\frac{a_{AU}}{d_{L}}$ such that $a$ is the semi-major axis in units of milli-arcseconds, $a_{AU}$ is the physical semi-major axis in units of AU and $d_{L}$ is the distance to the lens in units of kpc. In these models, $\boldsymbol{\mu}_{\boldsymbol{L},\sun}$ is treated as the proper motion of the binary lens system’s center of mass instead of the geometric midpoint of the binary lens. As such, we change the notation from $\boldsymbol{\mu}_{\boldsymbol{L},\sun}$ to $\boldsymbol{\mu}_{\boldsymbol{L,com},\sun}$ throughout this section. Furthermore, we amend the definition of $t_{0,\sun}$ for such binary-lens models, i.e, $t_{0,\sun}=t_{com,0,\sun}$, where $t_{com,0,\sun}$ is the time of closest approach between the lens center of mass and the source. The initial position of the lens system’s center of mass at $t_{com,0,\sun}$ is input as $\boldsymbol{X}_{\boldsymbol{L_{com},0},\sun}$.
@@ V.3 Full Keplerian Solutions
Using all eight Keplerian elements, the motion of the primary and secondary lenses around their center of mass ($\boldsymbol{X(t)}_{\boldsymbol{p},\boldsymbol{com},\sun}$ for primary and $\boldsymbol{X(t)}_{\boldsymbol{s},\boldsymbol{com},\sun}$ for secondary) at rest can be calculated. Once, the proper motion of the center of mass and the parallactic motion are taken into account, the primary and secondary lens trajectories are given by:
@@
We can simulate binary lens astrometry trajectories with Keplerian solutions implemented in BAGLE, presented in Figure 6. The code necessary to simulate the astrometric trajectories is displayed in Table 2. We use the PSBL parameterization PSBL_PhotAstrom_noPar_EllOrbs_Param1, while noting that BAGLE has multiple alternative parameterizations.
@@
3# Define model parameters and pass them into a model instance below.
4# mag_base and b_sff are arrays or lists. All other parameters are floats.
5# raL and decL must be defined when calling a parallax model.
6#root_tol is set by default to 1e-8. It is a parameter that can be passed to the model instance.
7
8psblorbits = model.PSBL_PhotAstrom_noPar_EllOrbs_Param1(
@@
10 beta, muL_E, muL_N, omega,
11 big_omega, i, e, tp, a, muS_E,
12 muS_N, dL, dS, b_sff, mag_src,
13 dmag_Lp_Ls, raL=None, decL=None,
14 root_tol=1e-8
@@
17# Define time array
18t = np.arange(t0 - 10*tE, t0 + 10*tE, 1)
19
20# Get resolved astrometry for lenses
21lens1, lens2 = psblorbits.get_resolved_lens_astrometry(t)
22
23# Get unlensed source trajectory
24source_unlensed = psblorbits.get_astrometry_unlensed(t)
25
26# Get resolved lensed images
27images_resolved = psblorbits.get_resolved_astrometry(t)
Table 2: Example bagle code used to initialize the PSBL orbital model and compute unlensed and lensed source astrometry.
V.4 Binary Lens Equation
@@
The complex conjugate of Equation  26 is:
@@
Differentiating Eqns. 26 and 31 gives
@@
We can then find the magnification of the source by the lens as:
@@
There are places in which $|J|\rightarrow 0$. This corresponds to infinite magnification (if the source were a point). The curves in the lens plane where this is true are “critical curves” and the corresponding curves in the source plane are called “caustics.” In the maps of magnification below (see §11), we show the source plane in which we can see caustics. We also show the caustics in Figures 13 and 14. When the source passes over a caustic, this is known as a “caustic crossing” and changes the number of lensed images of that source from 3 (when the source is outside of the caustic) to 5 (when it is inside the caustic).
@@
In order to avoid working with an equation with a mix of complex numbers and their conjugates, it is standard to plug Eqn. 31 into Eqn. 26 and simplify (see H. J. Witt, 1990; H. J. Witt & S. Mao, 1995). This yields a fifth-order polynomial known as the “lens polynomial.”
The lens polynomial is solved using numpy.roots() to obtain 5 solutions. The roots are then evaluated in the lens equation and confirmed to satisfy it. There may be some solutions that satisfy the lens polynomial that do not satisfy the lens equation. These are commonly called “ghost images” or “ghost roots” (e.g. V. Bozza et al., 2018). In addition to test the validity of a solution, BAGLE inputs a parameter root_tol for all binary lens models. root_tol is the tolerance threshold in comparing the polynomial roots to the physical solution. The root_tol defines how close a polynomial root must be to be accepted as a valid solution, and can be changed by the user. By default, it is set to to $10^{-8}$. In cases where the source is far away from the caustic, we select the correct 3 of the 5 solutions as those that satisfy the lens equation. All valid images contributing to the total magnification are retained, while non-physical roots are rejected using the Jacobian determinant criterion.
VI Binary Sources and Binary Lenses (BSBL)
Figure 8: Source and lens trajectories for a simulated binary-source, binary-lens microlensing event with circular orbits. The orange line is the primary source, the pink line is the secondary source, the black line is the primary lens, the purple line is the secondary lens, and the yellow lines are the image positions. The green line is the flux-weighted average of lensed source positions, as observed on the sky. Note that $mag_{S,pri}$ = 16 and $mag_{S,sec}$ = 17. The Einstein radius is $4.9$ mas. The simulation was run with a $t_{E}=231.15$ days with the following orbital parameters for the lens: $\omega_{pri}=30\degree$, $\Omega_{sec}=10\degree$, $\textit{i}=90\degree$, $e=0.2$, $\textit{P}=2722.46$ days, and the following orbital parmaeters for the source $\omega_{pri}=30\degree$, $\Omega_{sec}=10\degree$, $\textit{i}=90\degree$, $e=0.4$, $\textit{P}=6000$ days. The lenses have a mass of $m_{L,p}=$ 10 $M_{\odot}$ and $m_{L,s}=$ 8 $M_{\odot}$.
BAGLE can also simulate events with both binary sources and binary lenses. The orbital and lensing equations necessary to simulate a binary-source, binary-lens event in BAGLE are individually handled. This means that the equations necessary to simulate binary sources are from §IV, and the equations necessary to simulate binary lenses are from §V.
@@ VI Binary Sources and Binary Lenses (BSBL)
We present our astrometric simulations in Figure 8. BSBL_PhotAstrom_noPar_EllOrbs_Param2 is the BSBL model from BAGLE that is used to simulate an orbital solution for such a microlensing event. The BSBL model assumes that both - the source and the lens - display orbital motion. The figure includes the observed (unresolved) source position on the sky, i.e., the flux-weighted average of lensed source positions.
VII Validation of Models
Figure 9: Comparison of a simulated PSBL event with parallax between VBMicrolensing (using BinaryLightCurveParallax), MulensModel and BAGLE. The microlensing parameters in the SSB lens-frame are: $t_{0,\sun}=60478.49$ MJD, $u_{0,\sun}=1$, $t_{E,\sun}=100.4$ days, $q=0.3$, $\pi_{E}=[0.3,0.2]$. The magnification (top) and residuals with respect to the BAGLE model (bottom) are shown over time.
Figure 10: Comparison of a simulated BSPL event with parallax between VBMicrolensing, MulensModel and BAGLE. The microlensing parameters in the SSB lens-frame are: $t_{prim,0,\sun}=60478.49$ MJD, $t_{sec,0,\sun}=60476.09$ MJD, $u_{p,0,\sun}=1$, $u_{s,0,\sun}=1.22$, $t_{E,\sun}=100.4$ days, $q_{flux,s}=0.16$, $\pi_{E}=[0.3,0.2]$. The magnification (top) and residuals with respect to the BAGLE model (bottom) are shown over time.
@@ VII Validation of Models
In this section, we compare BAGLE with other contemporary microlensing models (VBMicrolensing and MulensModel) against simulated PSBL and BSPL events with parallax, but without orbital motion. Note that we present a comparison for point-source, point-lens models in J. Lu et al. (submitted).
A comparison with the inclusion of orbital motion and a comparison with BSBL is reserved for the future.
@@
We begin by creating a BAGLE model using the parameterization PSBL_Phot_Par_Param1. PSBL_Phot_Par_Param1 inputs quantities with reference to the geometric midpoint of the lens system, and calculates Earth’s position relative to the Solar System Barycenter over time. Event parameters in BAGLE are: $t_{0,\sun}=60478.49$ MJD, $u_{0,\sun}=1$, $t_{E,\sun}=100.4$ days, $q=0.3$, $\pi_{E}=[0.3,0.2]$. The event is located in the Galactic Bulge. The relative angle between the binary lens system and the $\mu_{\mathrm{rel},\sun}$ directional vector is $\phi=125\degree$. The two lenses have a separation of $\boldsymbol{s_{L}}(t_{0,\sun})=0.8$. In contrast to BAGLE, other packages prefer a geo-projected frame of reference (as described in J. Lu et al. (submitted)). The resulting geo-projected parameters are $t_{0,\earth_{r}}=60464.64$ MJD, $u_{0,\earth_{r}}=0.98$, $t_{E,\sun}=65.06$ days, $q=0.3$, $\pi_{E}=[-0.34,-0.13]$. After converting to the geo-projected frame of reference, an additional transformation must be applied to $\boldsymbol{u}_{\earth_{r}}$ and $t_{0,\earth_{r}}$ in order to switch from the geometric midpoint of the lens to the center of mass. The final values input to VBMicrolensing and MulensModel are $t_{0,\earth_{r}}=60460.41$ and $\boldsymbol{u}_{\earth_{r}}=0.78$. In the geo-projected frame, the heliocentric rectilinear frame has the same position and velocity as the true geocentric non-rectilinear frame at a reference time, denoted as $t_{0,par}$ in VBMicrolensing and MulensModel. For simplicity, we let $t_{0,par}=t_{0,\sun}=60478.49$ MJD in the comparison.
The comparison between VBMicrolensing, MulensModel, and BAGLE for PSBL events is shown in Figure 9. pyLIMA was not included because it uses VBMicrolensing as a backend. The difference in light curves between MulensModel and BAGLE is extremely small around  10-11. In contrast, differences of order 10-4 are apparent between BAGLE and VBMicrolensing. It is unclear why the difference between VBMicrolensing and BAGLE exists, and our future work involves resolving it.
@@
Next, a BSPL event pointing towards the Galactic Bulge was compared between VBMicrolensing, MulensModel and BAGLE. The parameterization BSPL_Phot_Par_Param1 was used to create a BAGLE model with the following microlensing parameters in the SSB reference frame: $t_{prim,0,\sun}=60478.49$ MJD, $t_{sec,0,\sun}=60476.09$ MJD, $u_{p,0,\sun}=1$, $u_{s,0,\sun}=1.25$, $t_{E,\sun}=100.4$ days, $q_{flux,s}=0.16$, $\pi_{E}=[0.3,0.2]$; where $u_{s,0,\sun}$ is the closest approach between the secondary source and lens, $t_{sec,0,\sun}$ is the time of this closest approach, and $q_{flux,s}$ is the ratio between the secondary source’s flux and the primary source’s flux. In the geo-projected frame, $t_{prim,0,\earth_{r}}=60464.64$ MJD, $t_{sec,0,\earth_{r}}=60461.01$ MJD, $u_{p,0,\earth_{r}}=0.98$, $u_{s,0,\earth_{r}}=1.22$, $t_{E,\earth_{r}}=65.06$ days, $\boldsymbol{\pi}_{E,\earth_{r}}=[-0.34,-0.13]$. The comparison between different BSPL models is presented in Figure 10. Differences of 10-4 are apparent between BAGLE and VBMicrolensing, likely due to differences in the parallax implementations. BAGLE and MulensModel have a much smaller difference around 10-15. This level of difference is consistent with the point-source, point-lens comparison (with parallax) presented in J. Lu et al. (submitted).
@@
Next, we compare the computational run time of BAGLE with VBMicrolensing and MulensModel. These runtimes are based on the time taken to model the events presented in Figures 9 and 10. Runtimes are calculated for models by generating mock data for an event with $t_{E}=100$ days sampling 2000 time steps over 5.5 years. The time tests are repeated 100 times and the mean and standard deviation are recorded for test. Tests were performed with BAGLE 1.0.5, VBMicrolensing v5.3.6, and MulensModel v3.3.1. A 2021 iMax equipped with a M1 processor using a Python 3.11 environment and the ipython kernel (Python gcc == 14.0.0) was used to conduct these tests. A summary of model-generation runtimes is presented in Table 3 and 4 for PSBL and BSPL events, respectively.
@@
The pre-instantiated runtime is applicable to running many different sets of model parameters for an event at the same sky location (i.e. model fitting). On the other hand, the full runtime is applicable to large-scale simulations of multiple events located at different sky coordinates. For pre-instantiated runtimes, we find that all packages are quick for BSPL events. BAGLE is the quickest for BSPL model generations when calculating the full runtime. However, BAGLE is the slowest for PSBL model generation (pre-instantiated and full). Future work involves implementing JAX (J. Bradbury et al., 2018) to improve the efficiency of the root solver algorithm in BAGLE.